An open container holds ice of mass 0.570 kg at a temperature of -18.0 ∘C . The mass of the container can be ignored. Heat is supplied to the container at the constant rate of 740 J/minute . The specific heat of ice to is 2100 J/kg⋅K and the heat of fusion for ice is 334×103J/kg.

PART A:

How much time tmelts passes before the ice starts to melt?

PART B:

From the time when the heating begins, how much time trise does it take before the temperature begins to rise above 0∘C?

PART C: How much power is radiated by a tungsten sphere (emissivity ϵ = 0.35) of radius 17 cm at a temperature of 19 ∘C?

PART D: If the sphere is enclosed in a room whose walls are kept at -5 ∘C, what is the net flow rate of energy out of the sphere?
Answer #1

heat required to change the temp. of ice from - 18 to 0

q1 = 0.57* 2100* 18 = 21546 J

heat required to chnage its state from ice to water

q2 = mL = 0.57* 334* 1000 = 190380 J

a)

now time it takes just before melting

t 1 = q1/ P = 29.11 mins

======

b)

t2 = (q1 + q2) /P = 286.386 mins

======

c) P = e * sigma * A T^4

P = 5.67* 10^-8* 0.35* 4* 3.14* 0.17^2* (273 + 19)^4

P = 52.368 W

======

d)

energy flow rate

E = 5.67* 10^-8* 0.35* 4* 3.14* 0.17^2* ( (273 + 19)^4 - ( 273 - 5)^4)

E = 15.208 J/s

========

Comment in case any doubt, will reply for sure.. Goodluck